X-ray with energy 12 keV is used to measure Bragg-reflection peaks. The third-order peak is measured for the X-ray beam incident at an angle of 30 degrees with

respect to the horizontal. What is the spacing between the atomic planes? What

is the angle at which the first-order reflection peak is observed?


1
Expert's answer
2021-11-19T10:51:32-0500

"2d\sin\alpha=n\lambda"


"h\nu=hc/\lambda=E\to \lambda=hc/E="


"=6.626\cdot10^{-24}\cdot3\cdot10^8/(12000\cdot1.6\cdot10^{-19})="


"=1.035\cdot10^{-10}\ (m)."


a. "2d\sin30°=3\cdot1.035\cdot10^{-10}\to d=3.1\cdot10^{-10}\ (m)"


b. "2\cdot3.1\cdot10^{-10}\cdot\sin\alpha=1\cdot1.035\cdot10^{-10}\to"


"\alpha=9.6°"